3. One analyzer has a resistor R assembled in series with a solenoid 1m long, has a voltage of 1000 and has a resistance RL = 6Ω.

Both ends of the circuit are connected to voltage U = 24V. It is known that the magnetic field generated by the electric current intensity is B = 4π x 10-4T. Let µo = 4π x 10-7SI.

   a. Calculate the current flowing through the circuit.

   b. Calculate the resistance R if the kinetic force of the analytical foot E = 4V.


1
Expert's answer
2020-09-28T08:07:49-0400

For solenoid,

Resistance R"_l" =6"\Omega"

length L=1m

Number of turns N=1000

Magnetic field B="4\pi\times 10^{-4}T"

"\mu_o=4\pi\times10^{-7}"

a) So Magnetic field for solenoid is given by the formula,

B="\frac{\mu_0Ni}{L}"


so i="\frac{BL}{\mu_oN}"

="\frac{4\pi\times10^{-4}\times 1}{4\pi\times10^{-7}\times1000}"


=1 A


b)Kinetic force at analytical foot E=4v

So using Ohm's law

v=ir

4=1"\times R"

R=4"\Omega"